OpenStudy (anonymous):

2 (radical sign) 3 * (radical sign) 12

OpenStudy (anonymous):

whats radical sign

OpenStudy (anonymous):

like square root

OpenStudy (saifoo.khan):

\[\Large 2\sqrt{3}*\sqrt{12}\] Like this?

OpenStudy (anonymous):

yea :)

OpenStudy (saifoo.khan):

12 is your answer! =)

OpenStudy (anonymous):

but how did u do it

OpenStudy (saifoo.khan):

\[2\sqrt{3}*2\sqrt{3}\]\[12\]

OpenStudy (anonymous):

\[2\sqrt{3}\cdot \sqrt{12} \]\[= 2\cdot \sqrt{3} \cdot \sqrt{12} \]\[= 2 \cdot\sqrt{3\cdot 12}\]\[=2\cdot \sqrt{36}\]\[=2\cdot \sqrt{6^2}\]\[= \ ?\]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (saifoo.khan):

uw.